What partial pressure of N2 gas (in mm Hg) is required to maintain a solubility of 4.17×10-3 g/L in water at 25 °C? kH for N2 at 25 °C is 6.47×10-4 mol/L·atm.
Answer #1

Answer:-

This question is solved by using the simple concept of Henry’s law of solubility of the gas in liquid which is directly proportional to the partial pressure of the gas over the solution.

The answer is given in the image,

Answer: z kya 6-47x10-4 mol/Laton Solubility = 4.14 x1d-3911 = 4•17x10 mol/l 28 = 1.49x10-4 moll solubility = KH. PM2 · PMy =